# EyeDentify: A Dataset for Pupil Diameter Estimation based on Webcam Images

### v1

@@ 3.4.2 Cropping the Eyes
The EAR for the left and right eye is defined as
$\text{EAR}_{\text{left}}=\frac{\|V_{\text{l1}}-V_{\text{l2}}\|}{\|H_{\text{l1}}-H_{\text{l2}}\|}\quad\text{and}\quad\text{EAR}_{\text{right}}=\frac{\|V_{\text{r1}}-V_{\text{r2}}\|}{\|H_{\text{r1}}-H_{\text{r2}}\|},$
@@
In humans, the complete blink process generally takes between 250 to 400 milliseconds [47], with the ”blackout duration,” when the upper lid covers the pupil, lasting from 40 to 200 milliseconds [11, 48, 47]. Given that our blink detection system identifies blinks based on the closing and opening of eyelids using EAR thresholds, we estimate a blink to last about 200 milliseconds. With a video capture rate of 30 frames per second, this translates to roughly 6 frames per blink. Detecting and excluding frames with blinks is crucial due to the frequency of blinks throughout the recording sessions. Figure 4 shows the pruning results for a particular participant across all sessions. The blue line indicates the total number of frames per session, while the orange line shows the frames retained after blink detection and removal.
Figure 4: Visualization of total frames vs. frames after pruning due to blink detection for one participant in all recording sessions (50 in total) as outlined in Section 3.4. Note that each recording takes a total of three seconds, which is why the impact of a blink and the amount of blinks can vary significantly (around 40 - 200 ms per blink).
Figure 5: Pupil diameter distribution of one participant during the recordings. A set of different pupil diameter measurements and webcam images were captured during the different three-second long sessions (in total, 50 sessions). The colors of the boxes indicate the display color used during the recordings (white, black, red, blue, yellow, green, gray, and white again).
4 Result
In this section, we describe the results of our dataset collection by describing the statistics of the participants (Section 4.1), the dataset distribution (Section 4.2), and first pupil diameter estimation results (Section 4.3). We use one common participant (number 6) to exemplify all visualizations in the following.
4.1 Dataset Statistics
In this study, we recruited 51 participants, including 39 males, 11 females, and 1 individual who preferred not to disclose their gender, all of whom consented to make their eye-cropped data public. The participants’ ages ranged from 21 to 44 years (M=27.58). They were predominantly workers or university students residing in Germany with origin nationalities in Eastern Europe, South Asia, Southeast Asia, North Africa, and North America. We obtained informed consent from all participants prior to the experiment to ensure compliance with the General Data Protection Regulation (GDPR). Additionally, participants were informed that they could opt out of the experiment at any time.
From an initial total of 226,912 image frames, our final dataset (after preprocessing described in Section 3.4) consists of 212,073 images of left and right eyes, curated from 51 participants after removing frames where blinks were detected. We meticulously organize each participant’s left and right eye images into separate directories. Additionally, we include a CSV file for each session. This file precisely documents the timestamp, session ID, gaze data, pupil data, and the path for each frame retained in that session, ensuring a comprehensive record of the eye-tracking data post-blink-filtering. This structured approach facilitates easy access and analysis of the dataset.
Figure 6: Two rug plot distributions of one participant with white display color (left) with pupil diameters ranging from 2.16mm to 2.28mm, and black display color (right) with pupil diameters ranging from 2.45mm to 2.80mm.
4.2 Dataset Distribution
Figure 5 shows the distribution of pupil diameters using a boxplot visualization. An in-depth look at the individual pupil diameters within one recording is shown in Figure 6 for white and black display colors. For the same participant, Figure 7 shows exemplary the corresponding webcam recording images for both eyes (cropped). As explained in Section 3.2, each session shows varied display colors. The ChameleonView web application captures a three-second webcam video recording for each session, synchronized with the pupil diameter data collected by the eye-tracker over the same duration. Since the eye-tracker records data at 90Hz, each session generates 270 data points. The mean and variance depicted in the boxplots are computed from these recordings.
Figure 7: Example webcam recording images (first and last frame) of both eyes of one participant (same as in Figure 6) with white display color (left) and black display color (right).
4.3 Pupil Diameter Estimation Result
Table 2 shows initial 5-fold cross-validation results on EyeDentify with the CNN models ResNet-18 and ResNet-50 [15]. For the left eye, the ResNet-18 model outperformed ResNet-50 with a lower MAE in both validation (0.0837 vs. 0.1001) and testing phases (0.1340 vs. 0.1426). Similarly, ResNet-18 again showed better performance for the right eye with a validation MAE of 0.1054 compared to ResNet-50’s 0.1089 and a test MAE of 0.1403 versus 0.1588 for ResNet-50. The results indicate that ResNet-18 consistently yields lower error rates than ResNet-50 across both eyes and both evaluation phases. The uncertainties represented by the standard deviations suggest some variability in the model performances across different subsets of the data, where ResNet-50 seems to be more consistent on the respective test partitions.
Eye Model Validation Test
MAE $\downarrow$ MAE $\downarrow$
Left ResNet-18 0.0837 $\pm$ 0.0135 0.1340 $\pm$ 0.0196
ResNet-50 0.1001 $\pm$ 0.0197 0.1426 $\pm$ 0.0167
Right ResNet-18 0.1054 $\pm$ 0.0173 0.1403 $\pm$ 0.0328
ResNet-50 0.1089 $\pm$ 0.0204 0.1588 $\pm$ 0.0203
Table 2: 5-fold cross-validation of ResNet-18 and ResNet-50 [15], evaluated separately for left and right eyes. Each group contains 10 randomly selected participants: 5 for validation and 5 for testing. The remaining participants were used to train the models. ResNet-18 performs the best for the pupil diameter estimation regarding mean values on the test partitions, whereas ResNet-50 shows a lower standard deviation, indicating more robustness for varied test partitions.
Figure 8 shows the class activation map (CAM) [56] of the last convolution layer of ResNet18 and ResNet50 respectively. The figure illustrates that ResNet18 primarily focuses on the external regions of the eyes, capturing environmental and skin color variations, with minimal attention (left eye) to no attention (right eye) given to the iris. For the left eye, ResNet50 also tends to focus strongly on the outer regions of the eye and mostly outside the iris, attempting to learn color intensities rather than concentrating on the iris or pupil. And, for the right eye, ResNet50 primarily focuses on the inner regions of the eye but neglects the surrounding color variations. These observations indicate that while focusing on the iris is crucial for accurately estimating pupil diameter, accounting for environmental variations can enhance performance and make the model more consistent.
Figure 8: ResNet18 ad ResNet50 test samples - CAM (class activation map) [56] visualizations with red display color. True and Predicted values show the original and estimated pupil diameter of the left and right eye in millimeters.
5 Limitations
EyeDentify substantially contributes to the eye monitoring research community. Nonetheless, it is important to acknowledge some limitations of our work to guide future research and application.
Firstly, the data was recorded using a single camera model, limiting the diversity in recording devices. Given that webcams vary widely in resolution and optical characteristics, we recommend future studies to validate and possibly enhance our initial pupil diameter estimation models across multiple camera types to ensure robustness.
Secondly, our EyeDentify primarily consists of participants without eyeglasses to maintain consistent conditions across all data. This exclusion limits the applicability of our findings to the general population, as eyeglass wearers comprise a significant demographic. Future work should consider this when they train models on EyeDentify for practical applications where users with eyeglasses are possible.
Lastly, EyeDentify does not account for personal eye characteristics that might influence measurements, such as eye color or health conditions. While we avoided collecting detailed personal meta eye data due to privacy concerns and the complexities of accurately self-reporting eye conditions, this omission can affect the applicability of our findings. Collaborating with ophthalmologists could enhance the collection process, ensuring accurate and privacy-sensitive inclusion of such data in future studies.
6 Future Work
Future work includes leveraging super-resolution models to enhance image quality, thereby providing more detailed features for deeper and more complex neural networks [32, 33]. For instance, face restoration or enhancement models such as GFPGAN [49] and CodeFormer [57], which have pre-learned facial features, can be particularly beneficial. Additionally, other state-of-the-art super-resolution models like regression-based transformed (SwinIR [27] or HAT [6]) or generative diffusion models (SR3 [42] or YODA [31]) can improve the resolution of webcam-captured eye images. These models can upscale images by factors of 2x, 3x, or 4x, facilitating the training of models on high-resolution images to estimate pupil diameter accurately, similar to gaze estimation using super-resolution [37, 52].
Furthermore, the features learned by models trained on high-resolution eye images for tasks such as gaze estimation (e.g., Appearance-based Gaze Estimation [7] and Super-Resolution for Appearance-Based Gaze Estimation [37]), blink detection (e.g., Real-Time Eye Gaze and Blink Estimation in Natural Environments [8]), or segmentation of eye parts and regions (e.g., Iris Segmentation Model [51]) can be employed for transfer learning. This approach can be applied to the readily available low-resolution images or the high-resolution images generated by super-resolution models, enhancing pupil detection’s overall performance and accuracy.
@@ 7 Societal Impact
We focus on publishing a dataset of everyday webcam images, which we hope will advance research about various fundamental physiological and psychological states and contribute to multiple human activity and healthcare domains. As such, these images and their associated pupil diameter information pose no immediate threat to individuals or organizations. The images were properly anonymized (by providing only cropped eyes). They do not depict any human or personal data that can be associated with any particular person due to the insufficient resolution of eye images for identification. As discussed in the limitations, EyeDentify may introduce biases inherent in the training data, i.e., missing nationalities or no eyeglasses. These biases could affect the quality and fairness of the pupil diameter estimations, particularly in diverse or underrepresented populations. Future work should aim to mitigate these biases and ensure that pupil diameter estimation technologies are inclusive.
@@ 8 Conclusion
We introduced EyeDentify, a novel dataset aimed at significantly enhancing eye monitoring research by enabling the development of models that estimate pupil diameter using standard webcam images. Such advancements are critical for understanding various physiological and psychological states through non-invasive methods, broadening the scope of applications in human-computer interaction, behavioral studies, and health diagnostics. Moreover, EyeDentify represents a significant advancement by addressing the scarcity of publicly available datasets that combine eye images with precise pupil diameter annotations. By focusing on standard webcam recordings, EyeDentify democratizes access to pupil-related research, potentially expanding its applicability in low-resource settings and everyday computing environments. Furthermore, our findings demonstrate that models trained on EyeDentify, particularly the ResNet-18 architecture, achieve promising results in estimating pupil diameters with a high degree of accuracy.

### v2

@@ 4.2 Cropping the Eyes
[…]
5 Model Training and Results
[…]
Instead of developing advanced deep learning models, our contribution and emphasis lies in providing publicly available dataset and the development of a practical web application for real-world pupil diameter estimation without specialized hardware. The best-performing ResNet models were integrated into our web application, PupilSense, and deployed on Hugging Face Spaces [^3] https://huggingface.co/spaces/vijulshah/pupilsense, enabling public access and advancing pupilometry research.
6 Web Application: PupilSense
We present a novel web application PupilSense shown in Figure 1. The application estimates pupil diameters from everyday images and videos. The application provides an in-depth analysis of the recordings, including Class Activation Maps (CAMs), which show the activated areas of the model influencing the output values based on the given input image, various graphs illustrating the predicted diameter values for the left and right pupils, Eye Aspect Ratios (EAR) in the event of blinking and a data frame table consolidating all the values in a single view for each frame in video inputs.
Within the application, users can upload photos or videos featuring either a person’s entire face or a close-up of the left or right eye. These images or videos can be captured using a smartphone or a webcam. The application allows users to adjust several settings, for instance, on uploading the media, users can select which pupil diameters to estimate, choosing from two primary options:
  • Both Pupils: This option automatically detects both the left and right eyes in the uploaded media, utilizing separate models to estimate the pupil diameters for each eye. The application identifies the face and crops out the eyes accordingly.
  • Single Pupil (Left or Right): Users can focus exclusively on the left or right eye. The application applies the corresponding model to estimate the diameter of the chosen pupil.
Users also have the option to choose between two model architectures: ResNet18 or ResNet50. Moreover, the application can detect blinks in the uploaded media if this feature is activated. This functionality employs an EAR threshold along with a Vision Transformer (ViT) model, analogous to the blink detection method described in Section 4. After uploading the media and configuring the desired settings, users can click the Predict Diameter and Compute CAM button to display results next to the uploaded media.
For images, the results include: (1) a cropped view of the left, right, or both eyes, depending on the user’s selection, (2) a CAM illustrates the areas activated in the model’s last convolutional layer based on the input image, and (3) the estimated pupil diameter for each eye (or the selected pupil).
For videos, results are displayed frame by frame, with each frame showcasing: (1) the cropped eye image, (2) the corresponding CAM image, and (3) the estimated pupil diameter for that specific frame.
As soon as the video starts processing, the resulting frames are played in a continuous loop at ten frames per second (Figure 1 - B) for easy viewing. For uploaded videos, an interactive line chart (Figure 1 - C) illustrating the estimated pupil diameter for the left or right pupil (or both) appears below the results. If blink detection is activated, the frames corresponding to blinking will not display estimated pupil diameters, resulting in gaps in the graph for those frames, as seen in (Figure 1 - C). When blink detection is enabled, each frame’s additional graph representing the EAR (Figure 1 - D) is displayed. This visual aid facilitates the identification of blinks, with marked horizontal green and red lines indicating the acceptable, rejectable, and gray zones for blink detection. Lastly, a data frame containing predicted diameters and EAR values is shown at the bottom in a Table format (Figure 1 - E), which can also be downloaded as a CSV file. It also contains two additional columns - one indicating the difference between the predicted diameter values for the corresponding frame and another indicating which one was greater.
The development of this app, along with its ability to visualize data and predictions, allows end users to gain a detailed understanding of the dynamics of pupil diameters. By providing transparency into how the models perform on images and videos, users can see the prediction process in action. This openness helps build trust in the application, as users can observe the models used, predictions made, and the steps the app takes to generate those predictions.
7 Limitations and Future Work
PupilSense currently supports only post-analysis, lacking real-time capabilities due to resource-limited hosting. More efficient models suitable for low-resource environments should be integrated to overcome this. Additionally, implementing real-time processing and estimation in the background while performing tasks related to specific medical diagnoses, such as ADHD, Alzheimer’s, Parkinson’s, or schizophrenia, is a potential area for future development. Additionally, validating models with diverse cameras, leveraging depth maps for estimations, and collaborating with ophthalmologists to improve data collection are essential. Mobile phone cameras should also be explored as a source for data.
The dataset’s reliance on a single camera model and exclusion of participants with eyeglasses or health conditions may impact its robustness. Additionally, the small size of pupil images limits feature extraction. Applying super-resolution (SR) techniques such as HAT [10] or SRResNet [34] and fine-tuning these models on eye-cropped datasets like FFHQ [27] or CelebA-HQ [23] could enhance detail, as shown in Shah et al. [52]. Combining SR with Pix2Net [24] for RGB-to-NIR image translation may further improve accuracy, especially in low-contrast conditions where distinguishing pupil features in darker irises is challenging.
@@ 8 Societal Impact
Our web application, PupilSense, prioritizes user privacy by not storing personal data. The dataset is released under the CC BY-NC 4.0 license, encouraging ethical, non-commercial use. Images and pupil diameter data are anonymized with low-resolution, cropped eye regions to prevent personal identification.
Despite these measures, biases in the training data—such as the lack of certain nationalities or individuals wearing eyeglasses—may impact the fairness and accuracy of estimations for underrepresented groups. Future research should aim to address these biases to enhance pupil diameter estimation systems, while also developing models that operate locally on users’ devices or transmit only cropped eye data to server-hosted models. This would further protect user privacy, ensure data sovereignty, and minimize the risks associated with the misuse of facial data.
@@ 9 Conclusion
In this work, we introduced PupilSense, a web application, along with a collected dataset aimed at advancing eye monitoring research by enabling the development of models that estimate pupil diameter using standard webcam images. Our dataset significantly contributes by addressing the shortage of publicly available datasets that provide eye images paired with precise pupil diameter annotations. Focusing on recordings from the webcams, our dataset opens up opportunities for pupil-related research in low-resource environments and everyday computing contexts. Our results demonstrate that models trained on our dataset, particularly the ResNet50 architecture, perform well in estimating pupil diameters. Additionally, we show the practicality of a web-based application for pupil diameter estimation that goes beyond controlled lab environments, offering an accessible solution for users without specialized technical knowledge, making it usable in natural, everyday settings.